# Modeling portfolio loss distribution under infectious defaults and immunization

### v3

@@ 1 Introduction
Modeling the joint loss distribution of a credit portfolio is a relevant and challenging task, fundamental for the pricing of credit derivatives. A vast literature that discusses the pricing of multiple-name credit derivatives (such as CDOs, i.e. Collateralized Debt Obbligations) flourished in the first decade of the 00s, period characterized by a strong expansion of such markets. Not surprisingly, the interest from the financial community declined after the Global Financial Crisis of 2007-08 due to the contraction of the CDOs markets, whose opacity and misleading pricing were considered among the main causes. Despite the reduced hype, and the tarnished reputation of CDO markets, synthetic CDOs contracts written on credit indices such as the CDX and iTraxx are still a relevant and growing market. These instruments allow investors to identify opportunities and hedging strategies. Regulators may also be interested in the information content of synthetic CDO tranches, as they embed the market estimates of default correlation and on the probability of extreme credit events (see e.g. ECB,, 2006; Scheicher,, 2008; Wojtowicz,, 2014; Gourieroux et al.,, 2021), providing market-based information relevant for the estimation of systemic risk (see e.g. Montagna et al.,, 2020; Gourieroux et al.,, 2021).
The primary objective of this paper is to develop a novel and tractable model for constructing the loss distribution of a credit portfolio. A secondary objective is to design and empirically validate a recursive algorithm for deriving this distribution. Our approach extends the contagion-based framework of Davis and Lo, (2001), but differs in key aspects. Instead of modeling pairwise contagion shocks, we assume that the default of a single name has systemic relevance, affecting the entire portfolio. We introduce an immunization mechanism that mitigates contagion effects and prevents feedback loops, simplifying calibration and improving computational efficiency. The model also relaxes the homogeneous portfolio assumption, allowing heterogeneous default probabilities and infection rates. To enhance flexibility, we propose two specifications integrating contagion with a common factor: (i) contagion conditional on a Gaussian factor, and (ii) a mixture distribution. These formulations produce tractable loss distributions suitable for practical applications, with parsimonious parametrization. Finally, we validate our framework using synthetic CDO tranches on the iTraxx index, obtaining a good empirical fit and meaningful interpretation of the parameters.
@@ 2 Literature review
One of the main problems in credit modeling is the default clustering: it has been observed, especially during recessions, that defaults are not uniformly spaced in time but rather tend to concentrate over small periods of time (see e.g. Azizpour et al.,, 2018 presenting evidence of default clustering in the US market).
@@
Duffie et al., (2009) empirically test whether under doubly stochastic assumption the process of cumulative defaults can be modeled as time changed Poisson process. They used intensity estimated on US data of corporate defaults from 1979 to 2004 derived in Das et al., (2007). Their result point out that the proposed model is not able to explain properly the default correlation of the data. Similarly, Azizpour et al., (2018) find that the amount of default clustering cannot be explained by the exposure to observable and latent factors, showing strong evidence of contagion dynamics.
@@
Frey et al., (2001) show that the correlation structure between obligors is not able to capture the dependence completely. It is necessary to introduce more relevant information on the lower tail dependence. This is equivalent to exogenously specify the appropriate copula from which the amount of default contagion is derived. Yu, (2007) works with an intensity based model where the default intensities are driven also by past defaults history, in addition to exogenous factors. A special case of this model is the copula approach presented by Schönbucher and Schubert, (2001). Still, the direct modeling of the correlation between default intensity processes introduces a weak dependency resulting in a lower joint probability of defaults than expected for highly correlated entities (Jouanin et al.,, 2001). Other approaches focus on distribution with fat tails for the modeling of latent factors (see e.g. Kalemanova et al.,, 2007; Albrecher et al.,, 2007), or model the evolution of latent variables of a structural model using jump-diffusion processes (Bujok and Reisinger,, 2012).
@@
One valid alternative is to explicitly enrich the modeling framework with contagion or infection mechanisms; this increases the probability of observing extreme losses in the portfolio and can also account for default clustering. Adding this particular dependency structure introduces a looping mechanism that makes calibration problematic: the probability of default of each name can impact and is impacted by the probability of default of the others. Several attempts have been proposed in order to resolve the looping issue when adding such effects. Relative to contagion mechanisms, Jarrow and Yu, (2001) impose a hierarchical approach and suggest to separate the firms into two groups: primary names can only default idiosyncratically while names in the second group can also default because of infection starting from the first set of names. Their work generalizes reduced-form models by making default intensities depend on counterpart default. The primary/secondary separation has been applied by other authors too, for example Rösch and Winterfeldt, (2008): they start from a one-factor model where the number of defaults of primary names can affect the default probabilities of secondary names.
@@
In Neu and Kühn, (2004), firms can have either mutually supportive or competitive relationships between them; a default will hence decrease the default probability of competitors but increase the same quantity for firms that had positive inflows from the name in distress (for example suppliers/clients).
@@
Egloff et al., (2007) instead add micro-structural dependencies via a directed weighted graph and show how even well diversified portfolios carry significant credit risk when such inter-dependencies are accounted for. The open problem in their approach (as well as in other network based models) is the calibration of the weights that form the network.
@@
Giesecke and Weber, (2004) add contagion processes to more standard common factor approaches while Frey and Backhaus, (2010) apply a Markov-chain model with default contagion to the problem of dynamically hedge CDO products. Nowadays the attention on contagion mechanise is increasing with a particular attention to explicit modeling on the interconnections among obligors (see Torri et al.,, 2018).
@@
Our work draws inspiration from the seminal paper of Davis and Lo, (2001) where firms can default in two ways, either idiosyncratically or via infection from other defaulting names. Unfortunately, in their most generic (and therefore elegant) specification of the model, one has to rely on Monte Carlo simulations in order to obtain the portfolio loss distribution. This is not computationally efficient for large portfolios since in a portfolio of size $n$, the variables to consider are of order $n^{2}$. The solution proposed by the authors leads to closed form results for the loss distribution at the cost of quite strict assumptions: the portfolio considered has to be homogeneous with respect to the probability of idiosyncratic default, the infection rates and the losses given default (LGD). Sakata et al., (2007) extended Davis and Lo’s original model by assuming that idiosyncratic defaults might in fact be avoided with the help of non-defaulted names. In their model there is hence not only a contagion that causes more defaults but also a positive effect due to the intervention of other names, called recovery spillage, that prevents entities from defaulting. The main results they obtain are similar to the ones presented in Davis and Lo, (2001) in terms of both complexity and assumptions required (homogeneous portfolio).
@@
Cousin et al., (2013) use a multiple period model where defaults happening at time $t$ can still cause contagion later on. In addition, they consider the case where more than one infection is needed to cause a default by contagion and, from a theoretical point of view, they relax several of Davis and Lo, (2001) assumptions. Unfortunately, the most generic specification of the model is computationally very demanding and the numerical applications shown in Cousin et al., (2013) are based on assumptions that are in line with Davis and Lo, (2001); in particular, the authors require that the Bernoulli variables used are exchangeable.
@@ 3 The proposed infection model
As previously mentioned, our model draws inspiration from the seminal paper of Davis and Lo, (2001). Given a portfolio of $n$ entities, Davis and Lo, (2001) model the default event at time $t$ of entity $i$ via the Bernoulli variable $Z_{i}(t)$ that can either take value 0 (indicating survival) or 1 (default) and that is constructed according to the following equation
$Z_{i}(t)=X_{i}(t)+(1-X_{i}(t))\cdot\left[1-\prod_{i\neq j}\left(1-X_{j}(t)\cdot Y_{i,j}(t)\right)\right].$ (1)
The main assumption behind equation (1) is the existence of $n$ variables $X_{i}(t),i=1,\cdots,n$ and $n\cdot(n-1)$ variables $Y_{i,j}(t),i,j=1,\cdots,n,\quad j\neq i$. Variable $X_{i}(t)$ is responsible for idiosyncratic default of the $i$ name, while $Y_{i,j}(t)$ governs the possibility that name $i$ is infected by name $j$. The additional assumption taken is that the variables $X_{i}(t)$ and $Y_{i,j}(t)$ are i.i.d. according to a Bernoulli distribution.
@@
where we have postulated the existence of $3n$ mutually independent Bernoulli variables $X_{1},...,X_{n},V_{1},...,V_{n},U_{1},...,U_{n}$ compared of the n+n(n1)n+n(n-1) of Davis and Lo, (2001).
@@ 3.1 Portfolio loss distribution at time $t$
We present an algorithm for calculating $P\{L_{n}(t)=h\}$ for a given integer $h$. The algorithm is similar, in spirit, to the one presented by Andersen et al., (2003) for conditionally independent models. Andersen et al., (2003) showed that an efficient way of performing the convolution of the independent conditioned default probabilities is to construct the portfolio loss distribution by adding each name j with j=1,,nj=1,...,n one by one via a recursive relationship.
@@
where $L_{n}^{I}(t)$ and $L_{n}^{C}(t)$ represents the units of losses due to idiosyncratic effects and to contagion events, respectively. More precisely, let us define three building blocks
$\begin{array}[]{lr}L_{n}^{I}(t):=\sum_{i=1}^{n}d_{i}\cdot X_{i}(t)&\text{Idiosyncratic driven losses}\\ L_{n}^{C}(t):=\sum_{i=1}^{n}d_{i}\cdot(Z_{i}(t)-X_{i}(t))&\text{Contagion driven losses}\\ L_{n}^{P}(t):=\sum_{i=1}^{n}d_{i}\cdot[1-X_{i}(t)]\cdot[1-U_{i}(t)]&\text{Potential losses}\end{array}$
where $L_{n}^{P}(t)$, instead, represents the number of units of potential losses in a (so far) uncontaminated world where the names added are not in default but don’t have a defensive strategy. We point out that if contagion has happened in the system, $L_{n}^{P}(t)$ is equal to zero, since any name without a defensive strategy is affected by the infection.
Consequently the distribution of $L_{n}(t)$ can be described as:
@@
where the integer $\bar{l}$ is the maximum amount of losses in the system. We further partition the probability space on the basis of the indicator function $\mathbb{I}_{C}$} that is equal to one if there is at least one infection active and zero otherwise:
Discarding the events with zero probability we rewrite Equation (5) as follows:
$\begin{array}[]{ll}P\{L_{n}(t)=h\}=&\sum_{{k=0}}^{\bar{l}-h}\left[P\{L_{n}^{I}(t)=h,L_{n}^{C}(t)=0,L_{n}^{P}(t)=k,{{\mathbb{I}_{C}}=0}\}\right]\,+\\ \quad&\sum_{{k=0}}^{{h}}\left[P\{L_{n}^{I}(t)=k,L_{n}^{C}(t)=h-k,L_{n}^{P}(t)=0,{{\mathbb{I}_{C}}=1}\}\right],\end{array}$ (7)
@@
Proposition 2 allows to construct the loss distribution recursively, adding progressively one name at the time (that is, we first consider a system with a single asset, then with two assets, and so on up to $n$). Since the recursion does not depend on the order chosen when adding names, the final distribution of the portfolio losses defined by $\alpha_{n}(h,k)$ and $\beta_{n}(k,h-k)$ is not affected by the order of the companies. We also point out that if the portfolio is equally weighted (as in the case of synthetic CDOs), and the loss given default (LGD) are assumed to be uniform across names, we can set $d_{j}=1\;\forall j$, so that $\bar{l}=n$, and then rescale accordingly the support of the loss distribution.
In practice, assuming constant LGD, we set $d_{j}=1\;\forall j$ and we compute the coefficients $\alpha_{n}(h,k)$ and $\beta_{n}(h,k)$ for Equation 7 using Algorithms 1 and 2. Both algorithms have cubic space and time complexity, although many combinations are skipped. The implementation is practical for applications with a moderate number of assets (up to $200$, as in the case of synthetic CDOs such as the iTraxx, where $n=125$), but for larger portfolio the algorithm becomes impractical. In such cases, a Monte Monte Carlo approach may be suitable, as discussed in Appendix B.
@@ 4.1 Conditional contagion loss distribution
In the first approach the defaults are driven by a common latent factor, and contagion is applied on top of the effect of the factor. We consider a credit portfolio in which default risk is modeled through a structural Vasicek-type framework (Vasicek,, 2002). For each obligor $i$, the default occurs between time 0 and $t$ when a latent variable $Z_{i}$ falls below a fixed threshold $\theta_{i}$. The latent variable follows the standard one-factor Gaussian model:

### v4

@@ 1 Introduction
Modeling the joint loss distribution of a credit portfolio is a relevant and challenging task, fundamental for the pricing of credit derivatives. A vast literature that discusses the pricing of multiple-name credit derivatives (such as CDOs, i.e. Collateralized Debt Obbligations) flourished in the first decade of the 00s, period characterized by a strong expansion of such markets. Not surprisingly, the interest from the financial community declined after the Global Financial Crisis of 2007-08 due to the contraction of the CDOs markets, whose opacity and misleading pricing were considered among the main causes. Despite the reduced hype, and the tarnished reputation of CDO markets, synthetic CDOs contracts written on credit indices such as the CDX and iTraxx are still a relevant and growing market. These instruments allow investors to identify opportunities and hedging strategies. Regulators may also be interested in the information content of synthetic CDO tranches, as they embed the market estimates of default correlation and on the probability of extreme credit events (see e.g. ECB, 2006; Scheicher, 2008; Wojtowicz, 2014; Gourieroux et al., 2021), providing market-based information relevant for the estimation of systemic risk (see e.g. Montagna et al., 2020; Gourieroux et al., 2021).
The primary contribution of this paper is to develop a novel and tractable model for constructing the loss distribution of a credit portfolio, providing also an analytical characterization of pairwise distributions among names. A secondary contribution is to design and empirically validate a recursive algorithm for deriving this distribution suitable for real-world asset pricing applications. Our approach extends the contagion-based framework of Davis and Lo (2001), but differs in key aspects. Instead of modeling pairwise contagion shocks, we assume that the default of a single name has systemic relevance, affecting the entire portfolio. We introduce an immunization mechanism that mitigates contagion effects and prevents feedback loops, simplifying calibration and improving computational efficiency. The model also relaxes the homogeneous portfolio assumption, allowing heterogeneous default probabilities and infection rates. To enhance flexibility, we propose two specifications integrating contagion with a common factor: (i) contagion conditional on a Gaussian factor, and (ii) a mixture distribution. These formulations produce tractable loss distributions suitable for practical applications, with parsimonious parametrization. Finally, we validate our framework using synthetic CDO tranches on the iTraxx index, obtaining a good empirical fit and meaningful interpretation of the parameters.
@@ 2 Literature review
One of the main problems in credit modeling is the default clustering: it has been observed, especially during recessions, that defaults are not uniformly spaced in time but rather tend to concentrate over small periods of time (see e.g. Azizpour et al., 2018 presenting evidence of default clustering in the US market).
@@
Duffie et al. (2009) empirically test whether under doubly stochastic assumption the process of cumulative defaults can be modeled as time changed Poisson process. They used intensity estimated on US data of corporate defaults from 1979 to 2004 derived in Das et al. (2007). Their result point out that the proposed model is not able to explain properly the default correlation of the data. Similarly, Azizpour et al. (2018) find that the amount of default clustering cannot be explained by the exposure to observable and latent factors, showing strong evidence of contagion dynamics.
@@
Frey et al. (2001) show that the correlation structure between obligors is not able to capture the dependence completely. It is necessary to introduce more relevant information on the lower tail dependence. This is equivalent to exogenously specify the appropriate copula from which the amount of default contagion is derived. Yu (2007) works with an intensity based model where the default intensities are driven also by past defaults history, in addition to exogenous factors. A special case of this model is the copula approach presented by Schönbucher and Schubert (2001). Still, the direct modeling of the correlation between default intensity processes introduces a weak dependency resulting in a lower joint probability of defaults than expected for highly correlated entities (Jouanin et al., 2001). Other approaches focus on distribution with fat tails for the modeling of latent factors (see e.g. Kalemanova et al., 2007; Albrecher et al., 2007), or model the evolution of latent variables of a structural model using jump-diffusion processes (Bujok and Reisinger, 2012).
@@
One valid alternative is to explicitly enrich the modeling framework with contagion or infection mechanisms; this increases the probability of observing extreme losses in the portfolio and can also account for default clustering. Adding this particular dependency structure introduces a looping mechanism that makes calibration problematic: the probability of default of each name can impact and is impacted by the probability of default of the others. Several attempts have been proposed in order to resolve the looping issue when adding such effects. Relative to contagion mechanisms, Jarrow and Yu (2001) impose a hierarchical approach and suggest to separate the firms into two groups: primary names can only default idiosyncratically while names in the second group can also default because of infection starting from the first set of names. Their work generalizes reduced-form models by making default intensities depend on counterpart default. The primary/secondary separation has been applied by other authors too, for example Rösch and Winterfeldt (2008): they start from a one-factor model where the number of defaults of primary names can affect the default probabilities of secondary names.
@@
In Neu and Kühn (2004), firms can have either mutually supportive or competitive relationships between them; a default will hence decrease the default probability of competitors but increase the same quantity for firms that had positive inflows from the name in distress (for example suppliers/clients).
@@
Egloff et al. (2007) instead add micro-structural dependencies via a directed weighted graph and show how even well diversified portfolios carry significant credit risk when such inter-dependencies are accounted for. The open problem in their approach (as well as in other network based models) is the calibration of the weights that form the network.
@@
Giesecke and Weber (2004) add contagion processes to more standard common factor approaches while Frey and Backhaus (2010) apply a Markov-chain model with default contagion to the problem of dynamically hedge CDO products. Nowadays the attention on contagion mechanise is increasing with a particular attention to explicit modeling on the interconnections among obligors (see Torri et al., 2018).
@@
Our work draws inspiration from the seminal paper of Davis and Lo (2001) where firms can default in two ways, either idiosyncratically or via infection from other defaulting names. Unfortunately, in their most generic (and therefore elegant) specification of the model, one has to rely on Monte Carlo simulations in order to obtain the portfolio loss distribution. This is not computationally efficient for large portfolios since in a portfolio of size $n$, the variables to consider are of order $n^{2}$. The solution proposed by the authors leads to closed form results for the loss distribution at the cost of quite strict assumptions: the portfolio considered has to be homogeneous with respect to the probability of idiosyncratic default, the infection rates and the losses given default (LGD). Sakata et al. (2007) extended Davis and Lo’s original model by assuming that idiosyncratic defaults might in fact be avoided with the help of non-defaulted names. In their model there is hence not only a contagion that causes more defaults but also a positive effect due to the intervention of other names, called recovery spillage, that prevents entities from defaulting. The main results they obtain are similar to the ones presented in Davis and Lo (2001) in terms of both complexity and assumptions required (homogeneous portfolio).
@@
Cousin et al. (2013) use a multiple period model where defaults happening at time $t$ can still cause contagion later on. In addition, they consider the case where more than one infection is needed to cause a default by contagion and, from a theoretical point of view, they relax several of Davis and Lo (2001) assumptions. Unfortunately, the most generic specification of the model is computationally very demanding and the numerical applications shown in Cousin et al. (2013) are based on assumptions that are in line with Davis and Lo (2001); in particular, the authors require that the Bernoulli variables used are exchangeable.
@@ 3 The proposed infection model
As previously mentioned, our model draws inspiration from the seminal paper of Davis and Lo (2001). Given a portfolio of $n$ entities, Davis and Lo (2001) model the default event at time $t$ of entity $i$ via the Bernoulli variable $Z_{i}(t)$ that can either take value 0 (indicating survival) or 1 (default) and that is constructed according to the following equation
$Z_{i}(t)=X_{i}(t)+(1-X_{i}(t))\cdot\left[1-\prod_{j\neq i}\left(1-X_{j}(t)\cdot Y_{i,j}(t)\right)\right].$ (1)
The main assumption behind equation (1) is the existence of $n$ variables $X_{i}(t),i=1,\cdots,n$ and $n\cdot(n-1)$ variables $Y_{i,j}(t),\;i,j=1,\cdots,n,\;j\neq i$. Variable $X_{i}(t)$ is responsible for idiosyncratic default of the $i$ name, while $Y_{i,j}(t)$ governs the possibility that name $i$ is infected by name $j$. The additional assumption taken is that the variables $X_{i}(t)$ and $Y_{i,j}(t)$ are i.i.d. according to a Bernoulli distribution.
@@
where we have postulated the existence of $3n$ mutually independent Bernoulli variables $X_{1},...,X_{n},V_{1},...,V_{n},U_{1},...,U_{n}$ compared of the n+n(n1)n+n(n-1) of Davis and Lo (2001).
@@ 3.1 Portfolio loss distribution at time $t$
We present an algorithm for calculating $P\{L_{n}(t)=h\}$ for a given integer $h$. The algorithm is similar, in spirit, to the one presented by Andersen et al. (2003) for conditionally independent models. Andersen et al. (2003) showed that an efficient way of performing the convolution of the independent conditioned default probabilities is to construct the portfolio loss distribution by adding each name $j$ with j=1,,nj=1,...,n one by one via a recursive relationship.
@@
where $L_{n}^{I}(t)$ and $L_{n}^{C}(t)$ represents the units of losses due to idiosyncratic effects and to contagion events, respectively. Beyond these two components, we consider a third group: firms that, in an uncontaminated world, have neither activated defensive strategies nor defaulted idiosyncratically. These firms remain solvent only because contagion has not yet occurred; in the presence of contagion, they would instead default. We therefore introduce a third quantity $L_{n}^{P}(t)$ which denotes the number of units of potential losses in a (so far) uncontaminated environment, corresponding to firms that are not in default but have not adopted any defensive strategy. Formally, these three building blocks are defined as:
$\begin{array}[]{lr}L_{n}^{I}(t):=\sum_{i=1}^{n}d_{i}\cdot X_{i}(t)&\text{Idiosyncratic driven losses}\\ L_{n}^{C}(t):=\sum_{i=1}^{n}d_{i}\cdot(Z_{i}(t)-X_{i}(t))&\text{Contagion driven losses}\\ L_{n}^{P}(t):=(1-{\mathbb{I}_{C}})\cdot\sum_{i=1}^{n}d_{i}\cdot[1-X_{i}(t)]\cdot[1-U_{i}(t)]&\text{Potential losses}\end{array}$
where $\mathbb{I}_{C}$} is an indicator function that is equal to one if there is at least one infection active and zero otherwise:
We underline that if contagion has happened in the system (i.e. ${\mathbb{I}_{C}}=1$), $L_{n}^{P}(t)$ is equal to zero, since any name without a defensive strategy is affected by the infection and thus they are counted in the $L_{n}^{C}(t)$ component. In contrast, $L_{n}^{P}(t)=0$ does not imply that there are no active contagions, as it can be that a company $j$ transmitted contagion, but any other company $i$ either defaulted idiosyncratically ($X_{i}=1$), or has immunity (i.e. $U_{i}=1$).
The distribution of $L_{n}(t)$ can be described as:
@@
where the integer $\bar{\ell}=\sum_{n}d_{i}$ is the maximum amount of losses in the system. We further partition the probability space on the basis of the indicator function $\mathbb{I}_{C}$}, and discarding the events with zero probability we rewrite Equation (6) as follows:
$\begin{array}[]{ll}P\{L_{n}(t)=h\}=&\sum_{{k=0}}^{\bar{\ell}-h}\left[P\{L_{n}^{I}(t)=h,L_{n}^{C}(t)=0,L_{n}^{P}(t)=k,{{\mathbb{I}_{C}}=0}\}\right]\,+\\ \quad&\sum_{{k=0}}^{{h}}\left[P\{L_{n}^{I}(t)=k,L_{n}^{C}(t)=h-k,L_{n}^{P}(t)=0,{{\mathbb{I}_{C}}=1}\}\right],\end{array}$ (7)
@@
Proposition 2 allows to construct the loss distribution recursively, adding progressively one name at the time (that is, we first consider a system with a single asset, then with two assets, and so on up to $n$). Since the recursion does not depend on the order chosen when adding names, the final distribution of the portfolio losses defined by $\alpha_{n}(h,k)$ and $\beta_{n}(k,h-k)$ is not affected by the order of the companies. We also point out that if the portfolio is equally weighted (as in the case of synthetic CDOs), and the loss given default (LGD) are assumed to be uniform across names, we can set $d_{j}=1\;\forall j$, so that $\bar{\ell}=n$, and then rescale accordingly the support of the loss distribution.
In practice, assuming constant LGD, we set $d_{j}=1\;\forall j$ and we compute the coefficients $\alpha_{n}(h,k)$ and $\beta_{n}(h,k)$ for Equation 7 using Algorithms 1 and 2. Both algorithms have cubic space and time complexity, although many combinations are skipped. The implementation is practical for applications with a moderate number of assets (up to $200$, as in the case of synthetic CDOs such as the iTraxx, where $n=125$), but for larger portfolio the algorithm becomes impractical. In such cases, a Monte Carlo approach may be suitable, as discussed in Appendix B.
@@ 4.1 Conditional contagion loss distribution
In the first approach the defaults are driven by a common latent factor, and contagion is applied on top of the effect of the factor. We consider a credit portfolio in which default risk is modeled through a structural Vasicek-type framework (Vasicek, 2002). For each obligor $i$, the default occurs between time 0 and $t$ when a latent variable $Z_{i}$ falls below a fixed threshold $\theta_{i}$. The latent variable follows the standard one-factor Gaussian model: